OpenStudy (anonymous):

Write an equation with the variable on each both side of the equal sign, at least one fractional cofficient, and a solution of-6.

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

i can think of this equation: (1/2) x +1=(2/3) x+2 verify whether -6 is the solution.

hartnn (hartnn):

i just put numbers randomly, like (1/2)x+1=bx+2 then found out b=2/3 by putting x=-6

OpenStudy (anonymous):

What is the answer then?

hartnn (hartnn):

(1/2) x +1=(2/3) x+2

OpenStudy (anonymous):

ok thank you

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

The distance traveled by a moving object is given by the formula: d=rt; where r is speed and t is time. Find the distance, d, if r-60mph and t=3.5hours.

hartnn (hartnn):

d=rt, so d=60*3.5 = ?

OpenStudy (anonymous):

what?